Brianna's family spent $134 on 2 adult tickets and 3 youth tickets at an amusement park. Max's family spent $146 on 3 adult tickets and 2 youth tickets. What is the price of a youth ticket?
Answer by mananth(16946)

x= cost of adult ticket
y= cost of youth ticket

2 x + 3 y = 134 .............1

3 x + 2 y = 146 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 3
-4 x -6 y = -268
9 x + 6 y = 438
Add the two equations
5 x = 170
/ 5
x = 34
plug value of x in (1)
2 x + 3 y = 134
68 + 3 y = 134
3 y = 134 -68
3 y = 66
y = 22 1
x= 34 cost of adult ticket
y= 22 cost of youth ticket
m.ananth@hotmail.ca